OpenStudy (landon34):

Find the GCF of n^3 t^2 and nt^4. I need help.

OpenStudy (whpalmer4):

GCF = greatest common factor We can find this by writing out the factoring of each quantity and comparing them \[n^3t^2 = n*n*n*t*t\]\[nt^4 = n*t*t*t*t\] Now, what is the largest number of \(n\)s that can be found in both factorings? Similarly, what is the largest number of \(t\)s that can be found in both factorings? Multiply those together and you have the GCF. To clarify, if you decide that the largest number of \(n\)s is 4, you would multiply \(n^4\), not just a \(4\)...

OpenStudy (landon34):

@whpalmer4

OpenStudy (landon34):

tell me when of if you get back

OpenStudy (whpalmer4):

I'm back...

OpenStudy (landon34):

ok

OpenStudy (whpalmer4):

So, what is the answer?

OpenStudy (landon34):

lemme read over your explanation. then Ill try to solve.

OpenStudy (whpalmer4):

seems like you could have done that without me being here...

OpenStudy (landon34):

ok I found out, thank you.

OpenStudy (whpalmer4):

What is the answer?

OpenStudy (whpalmer4):

@Landon34 if you want help, come back

OpenStudy (landon34):

ok.

OpenStudy (landon34):

yeah, Im only in 8th grade, and I hate algebra and pre-algebra. Im failing, and Im trying my best.

OpenStudy (landon34):

ill get one of my friends on here. @just_one_last_goodbye

OpenStudy (landon34):

wait, never mind I closed it, didnt even realize it. I'll make a new post.